OpenStudy (anonymous):

how do you find the interval for theta for which the graph is traced only once

OpenStudy (aum):

One full cycle for sin(x) is: \(0 \le x \le 2\pi\). One full cycle for sin(2x) is: \(0 \le x \le \large \frac{2\pi}{2}=\pi\). One full cycle for sin(3x) is: \(0 \le x \le \large \frac{2\pi}{3}\). One full cycle for sin(k*x) is: \(0 \le x \le \large \frac{2\pi}{k}\).

OpenStudy (sleepyjess):

\(\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\\\huge\cal\color{red}W\color{orange}E\color{goldenrod}L\color{yellow}C\color{lightgreen}O\color{darkgreen}M\color{turquoise}E~~\color{royalblue}T\color{purple}O~~\color{#00bfff}{Open}\color{#11c520}{Study}\\\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\)